Lucy bought a car for $23,000. The value of the car depreciates at about 17% each year. When will the value of the car be $10000?
Answer by stanbon(75887)

Lucy bought a car for $23,000. The value of the car depreciates at about 17% each year. When will the value of the car be $10000?
------
Each year the car is worth 83% of its previous year's worth.
-------------------
Solve 10,000 = 23000(0.83)^k
----
(0.83)^k = 10/23
k = [log(10/23)]/[log(0.83)] =
k = 4.47 years
====================
Cheers,
Stan H.
====================